# Navigating the Landscape of Hint Generation Research: From the Past to the Future

### v1

@@ 5 Conclusion
In this paper, we consolidate prior research in hint generation, bridging the gap between research in education and cognitive science, and research in AI and natural language processing. Based on our findings, we propose a roadmap for the future research in hint generation, where we provide a rethinking of the formal task definition, a brief review of research areas that can inform the design of future systems, open challenges for effective hint generation systems, and the ethical considerations. Although hint generation has a long history dating back over three decades Hume et al. (1996), recent advances in natural language processing could serve as useful for future hint generation systems. Beyond education, hint generation is also an excellent atomic task to measure a system’s ability to personalize content to user needs and requirements. We invite researchers to foster a community, develop new benchmarks, create shared tasks and workshops for automatic hint generation.
@@
Due to the rich and diverse literature on intelligent tutoring systems, we limit our survey to research directly relevant to hint generation and do not cover other types of feedback.
Acknowledgements
We would like to thank Shivani Kapania for identifying ethical considerations and proof-reading this work.

### v2

@@ 5 Conclusion
In this paper, we consolidate prior research in hint generation, bridging the gap between research in education and cognitive science, and research in AI and natural language processing. Based on our findings, we propose a roadmap for the future research in hint generation, where we provide a rethinking of the formal task definition, a brief review of research areas that can inform the design of future systems, open challenges for effective hint generation systems, and the ethical considerations. Although hint generation has a long history dating back over three decades Hume et al. (1996), recent advances in natural language processing could serve useful for future hint generation systems. Beyond education, hint generation is also an excellent atomic task to measure a system’s ability to personalize content to user needs and requirements. We invite researchers to foster a community, develop new benchmarks, create shared tasks and workshops for automatic hint generation.
@@
Due to the rich and diverse literature on intelligent tutoring systems, we limit our survey to research directly relevant to hint generation and do not cover other types of learning-related feedback.